Q1.

A labor force survey for UG Ll 00 economy suggests that the total number of employed are

11 million, with an unemployment rate of 5.5 %. If the labor force participation rate is 62%,

calculate

i. Unemployment rate

ii. Labor Force

iii. Assume that there are 20,000 discouraged workers, how will the unemployment rate

change.

iv.

Assume instead, there are 20,000 dishonest workers, how will the unemployment rate

change.



1
Expert's answer
2021-06-29T03:40:02-0400

i) Unemployment rate

Employment rate: 100 - 5.5 = 94.5%

94.5% = 11million

5.5% ?

(5.5 * 11m)/94.5%

= 640,212


ii) Labor force

Labor force = number of persons employed + number of persons unemployed

= 11 million + 640,212

= 11,640,212


iii) Unemployment rate is based solely on the active labor force number and therefore does not take into account the number of discouraged workers in a country. 20,000 discouraged workers will reduce the unemployment rate to:

640,212 - 20,000 = 620,212 (new unemployment rate number)

620,212/11,620,212 * 100 = 5.3%

Unemployment reduces from 5.5% to 5.3%


iv)Dishonest workers are likely to be dismissed from their jobs thus increasing the unemployment rate.

640,212 + 20,000 = 660,212 (new unemployment rate number)

660,212/11,660,212 * 100 = 5.7%

Unemployment increases from 5.5% to 5.7%